# Metacognition and Uncertainty Communication in Humans and Large Language Models

### v1

@@ Abstract
Metacognition—the capacity to monitor and evaluate one’s own knowledge and performance—is foundational to human decision-making, learning, and communication. As large language models (LLMs) become increasingly embedded in high-stakes decision contexts, it is critical to assess whether, how, and to what extent they exhibit metacognitive abilities. Here, we provide an overview of current knowledge of LLMs’ metacognitive capacities, how they might be studied, and how they relate to our knowledge of metacognition in humans. We show that while humans and LLMs can sometimes appear quite aligned in their metacognitive capacities and behaviors, it is clear many differences remain; attending to these differences is crucial not only for enhancing human-AI collaboration, but also for promoting the development of more capable and trustworthy artificial systems. Finally, we discuss how endowing future LLMs with more sensitive and more calibrated metacognition may also help them develop new capacities such as more efficient learning, self-direction, and curiosity.
@@ 1 Introduction
Metacognition refers to the human capacity to monitor, assess, and regulate our own cognitive processes and mental states. Within the broader field of metacognition, confidence judgments—subjective evaluations of the certainty of our own decisions—play a particularly important role (Fleming, 2024). Confidence judgments help individuals assess when their choices are likely correct or potentially erroneous. Expressing confidence also enables effective communication about uncertainty, facilitating collaboration by helping teams leverage collective knowledge and recognize trustworthy decisions (Frith, 2012).
@@
Recent developments in artificial intelligence (AI) have placed considerable attention on uncertainty and its effective communication to human users. Large language models (LLMs), in particular, increasingly serve in advisory roles, providing recommendations, explanations, and answers to diverse inquiries. Consequently, LLMs must be able communicate uncertainty effectively, enabling humans to appropriately calibrate their reliance on AI-generated recommendations and to understand clearly when such advice is dependable (Zhang et al., 2020; Steyvers et al., 2025; Steyvers & Kumar, 2024). Therefore, it is important to understand LLMs’ metacognitive capabilities, and to explore their capacity to communicate internal uncertainty, in order to facilitate their optimal collaboration with humans and future development potential.
@@
Here we examine key recent findings in LLMs’ metacognitive capabilities in relation to the human literature, highlighting the methods for evaluating confidence capacities and expressions of uncertainty with emphasis on human-LLM collaboration. Throughout, we provide insights into the parallels and divergences between human and LLM metacognition, illustrate contemporary challenges in AI uncertainty quantification, and discuss potential pathways for enhancing metacognitive interactions between humans and LLMs. We end by highlighting how improving metacognitive abilities in LLMs may facilitate their future development of other capacities associated with intelligence.
2 Metacognitive Capabilities of LLMs
@@
A key question regarding LLMs’ metacognition is whether they can accurately recognize and adequately communicate their own knowledge boundaries. Existing research is mixed in its conclusions. Some studies suggest that LLMs demonstrate limited metacognitive insight and struggle to recognize gaps in their own knowledge, leading to conclusions that LLMs lack essential metacognitive capabilities (Griot et al., 2025). Yet other findings suggest that LLMs can indeed detect their knowledge boundaries and can discriminate effectively between problems they can solve correctly and those for which they may fail (Yin et al., 2023; Kadavath et al., 2022; Steyvers et al., 2025). A contributing factor to these seemingly conflicting results is the large diversity in methods used to evaluate LLM confidence. Broadly, two approaches dominate current research: explicit confidence elicitation and implicit confidence estimation.
Explicit confidence elicitation involves prompting the model to verbalize its level of certainty—either through qualitative statements (e.g., “I’m not sure”) or quantitative confidence judgments expressed as percentages or probabilities (e.g., “I’m 70% sure”) (Cash et al., 2024; Griot et al., 2025). These outputs are generated via text, relying on the model’s ability to represent and articulate its own uncertainty in language. In contrast, implicit confidence estimation seeks to infer model confidence by measuring the likelihood assigned to tokens at the output layer of the LLM (Steyvers et al., 2025). For example, when answering a multiple-choice question with options A, B, C, and D, a confidence measure can be derived from the probability distribution over these choices. This method extends to open-ended questions through the p(true) approach (Kadavath et al., 2022), where the model first generates an answer and is then separately queried about its truthfulness. The confidence score is then calculated based on the probability that the model assigns to“true” versus “false” responses. These behaviors suggest the presence of metacognition through serving as implicit indicators, similar to how indirect evidence has been interpreted in non-human animal research: rats can “report” higher confidence in a decision by waiting longer for a food reward, and their behavioral patterns precisely map onto explicit confidence reports in humans and monkeys (Stolyarova et al., 2019). Studies consistently find that implicit confidence measures derived from token likelihoods tend to exhibit greater metacognitive sensitivity (see next section) compared to verbalized confidence elicited through explicit prompting (Xiong et al., 2024). This discrepancy highlights an important distinction between what models internally “know” (or represent) and what they explicitly express.
Additionally, metacognitive performance—by either of these approaches—is not uniform across different LLMs. Differences in model architecture and scale significantly impact the degree of apparent metacognitive capability. Larger and more sophisticated models tend to show improved calibration—meaning their expressed confidence better matches their actual accuracy—as well as heightened metacognitive sensitivity—meaning their confidence judgments can better differentiate between correct and incorrect responses (Kadavath et al., 2022; Xiong et al., 2024). These observations suggest that scaling up model size and optimizing architectural components may enhance LLMs’ metacognitive capacities. To fully comprehend and measure LLM metacognition, we must carefully think about evaluation methods and be aware of the inherent differences between models.
3 Methods to Assess Confidence-Accuracy Relationship
@@
Several metrics have been used to assess the relationship between confidence and accuracy across both humans and AI systems. While these metrics differ across disciplines, with some metrics originating in computer science and others in cognitive science, the metrics reveal two key facets of metacognition, briefly introduced above: metacognitive calibration and metacognitive sensitivity (Fleming, 2023; Lee et al., 2024; Li & Steyvers, 2025). For visual reference, Figure 1 illustrates both concepts, and compares them to empirical results for GPT-3.5 on a multiple-choice question-answering task. In this section, we describe these facets in more detail.
@@
Metacognitive sensitivity quantifies how “diagnostic” confidence judgments are of decisional accuracy—i.e., whether they reliably discriminate between correct or incorrect answers (Figure 1, top row). Within the human literature, metacognitive sensitivity metrics include phi ($\phi$) correlation (i.e., the correlation between accuracy and confidence across trials), the area under the type 2 receiver operating characteristic curve (AUC)—corresponding to the probability that a randomly sampled correctly answered question receives a higher confidence score than a randomly sampled incorrectly answered question—and a signal detection theoretic metric known as meta-d’ (analogous to d’ from signal detection theory), among others (Fleming & Lau, 2014). In contrast, metacognitive calibration refers to whether an observer reports a generally appropriate level of confidence given their probability of being correct. For example, if an individual—or an LLM—reports 75% confidence across multiple trials, calibration can be considered optimal when the actual proportion of correct answers in those trials is also 75% (Maniscalco et al., 2024). The Expected Calibration Error (ECE) is often used in computer science research to summarize the overall discrepancy between confidence and accuracy. ECE is typically computed by binning predictions according to confidence levels and comparing average confidence within each bin to the empirical accuracy. Calibration curves—graphs plotting model confidence against observed accuracy—are also commonly used to visualize calibration performance (Figure 1, bottom row). A perfectly calibrated system would exhibit a calibration line that falls on the diagonal (i.e., predicted confidence equals actual accuracy at all levels). Deviations from this line reflect systematic biases such as overconfidence (when predicted confidence exceeds accuracy) or underconfidence (when accuracy exceeds confidence). However, note that in the literature on human metacognition, apparent over- or under-confidence may in fact be mathematically optimal when considering reward functions or the observer’s global strategy or goals, such as whether it is more desirable to maximally avoid high-confidence errors given the consequences of such errors in the environment (Maniscalco et al., 2024).
@@
Figure 1: Demonstrations of confidence-accuracy relationships using cartoons and an empirical example based on empirical results from GPT-3.5 from (Steyvers et al., 2025), focusing on metacognitive sensitivity and calibration. Top row: Confidence distributions for correct (green) and incorrect (blue) answers allow assessment of metacognitive sensitivity. Illustrative results show examples of different degrees of separations between the distributions reflecting different degrees of metacognitive sensitivity. The empirical results using GPT-3.5 show modest separation, with the area under the curve (AUC = 0.778) reflecting the probability that a randomly selected correct answer is assigned higher confidence than a randomly selected incorrect answer. Bottom row: Metacognitive calibration can be seen by plotting accuracy as a function of confidence. The illustrative results show examples of over-confidence, under-confidence, and properly calibrated confidence (points directly along the diagonal). The empirical results for GPT-3.5 on a multiple-choice question-answering task show overconfidence, as the model’s predicted confidence exceeds observed accuracy.
@@ 4 Comparing Human and LLM Metacognitive Architecture and Behavior
There are several notable parallels between how humans and LLMs not only generate and calibrate confidence, but also express it. These similarities may seem surprising, given the fundamental architectural and cognitive differences between humans and LLMs, yet important differences also remain; exploring these differences and their consequences on collaborative behavior may be crucial to effective human-LLM collaboration.
@@ 4.1 Similarities Between Humans and LLMs
One point of convergence may be with some mechanisms thought to generate confidence. In LLMs, one approach to estimating confidence leverages their probabilistic nature: the model can be prompted multiple times with the same question, and confidence can be inferred from the consistency of responses (Xiong et al., 2024) — similar to the other implicit measures of metacognition discussed above — or their semantic coherence (Farquhar et al., 2024). Interestingly, this approach is similar to a proposed theoretical framework for human confidence, in which subjective certainty arises from the self-consistency of internally generated candidate answers (Koriat, 2012). Although developed independently in AI and cognitive psychology, both approaches suggest that consistency across internally simulated alternatives may serve as a basis for confidence.
Another similarity concerns the outwardly-visible behavioral patterns of calibration and sensitivity. Recent work has shown that when given the same task, LLMs and humans both tend to exhibit overconfidence, and both can achieve a similar degree of metacognitive sensitivity—that is, their confidence ratings are similarly diagnostic of accuracy (Cash et al., 2024). The tendency toward overconfidence has long been observed in human cognition (Kelly & Mandel, 2024), and appears to extend to large language models as well, possibly due to inductive biases or training data characteristics (Zhou et al., 2024).
@@
Another important parallel is that metacognitive abilities for both humans and LLMs can be improved through training. In the case of LLMs, research has shown that confidence verbalization (specifically, metacognitive sensitivity) can be improved through finetuning approaches that reward the LLM for accurately conveying uncertainty to a listener (Stengel-Eskin et al., 2024). Similarly, humans can improve metacognitive sensitivity and calibration through training (Carpenter et al., 2019). More specifically, providing feedback, encouraging reflective reasoning, and explicitly targeting cognitive biases can reduce human miscalibration of confidence (Kelly & Mandel, 2024). However, the way in which feedback is presented may change how an observer – biological or artificial – calibrates their confidence based on the strategy the observer is trying to optimize (Maniscalco et al., 2024).
Finally, metacognition in humans is thought to rely on introspection-like processes, defined specifically by the privileged access we have to our own thoughts over those of others (i.e., the difference between metacognition and theory of mind). Similarly, it has been suggested that LLMs can better predict their own behavior than the behavior of another LLM, which some researchers interpret to imply the presence of such privileged access in the LLMs tested (Binder et al., 2024). Evidence of introspective-like capacities may also come from LLMs’ demonstrated ability to describe their own behaviors even when they are not explicitly described in their training data, including ‘backdoors’ where models show unexpected or undesirable behaviors only under certain trigger conditions (Betley et al., 2025).

### v2

@@ Abstract
Metacognition—the capacity to monitor and evaluate one’s own knowledge and performance—is foundational to human decision-making, learning, and communication. As large language models (LLMs) become increasingly embedded in both high-stakes and widespread low-stakes contexts, it is important to assess whether, how, and to what extent they exhibit metacognitive abilities. Here, we provide an overview of current knowledge of LLMs’ metacognitive capacities, how they might be studied, and how they relate to our knowledge of metacognition in humans. We show that while humans and LLMs can sometimes appear quite aligned in their metacognitive capacities and behaviors, it is clear many differences remain; attending to these differences is important for enhancing human-AI collaboration. Finally, we discuss how endowing future LLMs with more sensitive and more calibrated metacognition may also help them develop new capacities such as more efficient learning, self-direction, and curiosity.
@@ 1 Introduction
Metacognition refers to the human capacity to monitor, assess, and regulate our own cognitive processes and mental states. It is foundational for learning, decision-making, and communication. Within this framework, confidence judgments and uncertainty representations play central roles. Confidence is a specific form of certainty and involves an explicit evaluation that a given choice is correct. Confidence is therefore tied directly to evaluating one’s own decision (Pouget et al., 2016). In contrast, uncertainty can be considered the broader internal representation of possible states or outcomes, which may or may not be explicitly expressed. Therefore, confidence is a particular, overt expression of uncertainty, and together these constructs provide measurable indicators of metacognition (Fleming, 2024; Pouget et al., 2016).
@@
Importantly, confidence does not only shape an individual’s own decisions but also serves a communicative function. Expressing confidence enables humans to coordinate effectively, by signaling when their judgments are likely trustworthy and when they may be error-prone (Frith, 2012). This communication of uncertainty allows groups to integrate knowledge efficiently and to calibrate trust across team members. Recent developments in artificial intelligence (AI) have placed considerable attention on uncertainty and its effective communication to human users. Large language models (LLMs), in particular, increasingly serve in advisory roles, providing recommendations, explanations, and answers to diverse inquiries. Consequently, LLMs must be able communicate uncertainty effectively, enabling humans to appropriately calibrate their reliance on AI-generated recommendations and to understand clearly when such advice is dependable (Steyvers et al., 2025b; Steyvers & Kumar, 2024). Therefore, it is important to understand LLMs’ metacognitive capabilities, and to explore their capacity to communicate uncertainty, in order to facilitate their effective use in human collaboration.
@@
Here we examine key recent findings in LLMs’ metacognitive capabilities in relation to the human literature, highlighting the methods for evaluating internal uncertainty and explicit confidence reporting with an emphasis on human-LLM collaboration. Throughout, we provide insights into the parallels and divergences between human and LLM metacognition, and discuss potential pathways for enhancing metacognitive interactions between humans and LLMs. In closing, we consider how advances in LLM metacognition might contribute to the emergence of other cognitive functions relevant to intelligence.
2 Confidence and Uncertainty Quantification in LLMs
@@
A key question regarding LLMs’ metacognition is whether they can accurately recognize and adequately communicate their own knowledge boundaries. Existing research is mixed in its conclusions. Some studies suggest that LLMs demonstrate limited metacognitive insight and struggle to recognize gaps in their own knowledge, leading to conclusions that LLMs lack essential metacognitive capabilities (Griot et al., 2025). Yet other findings suggest that LLMs can indeed detect their knowledge boundaries and can discriminate effectively between problems they can solve correctly and those for which they may fail (Kadavath et al., 2022; Steyvers et al., 2025b); see Figure 1 for a few examples. A contributing factor to these seemingly conflicting results is the diversity in methods used to quantify LLM uncertainty and the different ways in which the term confidence is used in the machine learning and psychology literature. Broadly, two approaches dominate current research: explicit and implicit methods to assess uncertainty.
Implicit methods seek to infer model uncertainty by either consistency-based methods or token likelihoods. With consistency-based methods, the agreement between multiple generated answers from an LLM determines uncertainty: If the model is certain, the same question tends to produce more consistent answers (Liu et al., 2025). With the token likelihood method, in contrast, the likelihood assigned to tokens at the output layer of the LLM is taken as a measure of uncertainty (Steyvers et al., 2025b; Liu et al., 2025). For example, when answering a multiple-choice question with options A, B, C, and D, the model generates a probability distribution over these choices that reflects its internal uncertainty about the answer option to generate. Unlike consistency-based methods, which often rely on sampling variability introduced through parameters such as temperature, the token likelihood approach uses the distribution computed during a single forward pass and does not depend on additional randomness or counterfactual generations. The token likelihood method extends to open-ended questions through the p(true) approach (Kadavath et al., 2022), where the model first generates an answer and is then prompted with a follow-up query such as “Is this statement true or false?”. The probability assigned to “true” versus “false” tokens is then taken as the confidence score. Although this approach involves issuing an additional query, it is still considered an implicit method because the model is not explicitly asked to verbalize its level of confidence; rather, researchers infer confidence from token likelihoods in the follow-up response.
These implicit measures of confidence can serve as indirect evidence for metacognitive computations, similar to how indirect evidence has been interpreted in non-human animal research: rats can “report” higher confidence in a decision by waiting longer for a food reward, and their behavioral patterns precisely map onto explicit confidence reports in humans and monkeys (Stolyarova et al., 2019). However, the true test for LLM metacognitive confidence is through explicit methods that involve prompting the model to verbalize its own level of confidence—either through qualitative statements (e.g., “I’m not sure”) or quantitative confidence judgments expressed as percentages or probabilities (e.g., “I’m 70% sure”) (Cash et al., 2025; Griot et al., 2025; Steyvers et al., 2025a)—rather than an external observer inferring the uncertainty present in the model. These outputs are generated via text, relying on the model’s ability to represent and articulate its own uncertainty in language.
Both implicit and explicit methods have been used by various groups to assess LLMs’ metacognitive performance, i.e. the degree to which LLMs’ confidence (or uncertainty) reflects their task accuracy. These studies find that differences in model architecture and scale can influence how well LLMs express confidence in ways reflect their underlying accuracy. For instance, some models appear better able to express high confidence for correct answers and lower confidence for incorrect ones (Kadavath et al., 2022; Xiong et al., 2024), or to express confidence levels that more closely match their actual probability of being correct. Yet direct comparisons between LLMs’ metacognitive capacities often involve mixed assessments, with some groups relying on explicit and others on implicit measures, and studies have consistently found that implicit confidence measures derived from token likelihoods tend to exhibit greater trial-by-trial correspondence between confidence and task accuracy than does verbalized confidence elicited through explicit prompting (Xiong et al., 2024). This discrepancy highlights an important distinction between what models internally “know” (or represent)—which can be accessed by an external observer—and what they can explicitly express. This underscores the need for consistent and precise evaluation methods to meaningfully assess metacognitive capabilities across LLMs.
3 Metrics for Assessing the Confidence-Accuracy Relationship
@@
Several metrics have been used to assess the relationship between confidence and accuracy across both humans and AI systems. While these metrics differ across disciplines, with some metrics originating in computer science and others in cognitive science, the metrics reveal two key facets of metacognitive ability: metacognitive sensitivity and metacognitive calibration (Fleming, 2023; Lee et al., 2025; Li & Steyvers, 2025). For visual reference, Figure 1 illustrates both concepts, and compares them to empirical results for GPT-3.5 on a multiple-choice question-answering task and GPT-4.1 on a short-answer trivia task.
Metacognitive sensitivity (also called metacognitive discrimination accuracy, relative accuracy, or monitoring resolution) quantifies how “diagnostic” confidence judgments are of decisional accuracy—i.e., whether they reliably discriminate between correct or incorrect answers (Figure 1, top row). Within the human literature, metacognitive sensitivity metrics include phi ($\phi$) correlation (i.e., the correlation between accuracy and confidence across trials), the area under the type 2 receiver operating characteristic curve (AUROC2)—corresponding to the probability that a randomly sampled correctly answered question receives a higher confidence score than a randomly sampled incorrectly answered question—and a signal detection theoretic metric known as meta-d’ (analogous to d’ from signal detection theory), among others (Fleming & Lau, 2014). Worth noting here is that most measures of metacognitive sensitivity (with the exception of meta-d’, of those discussed here) are ‘contaminated’ by type 1 accuracy, or the observer’s capacity to complete the target task. This means that an apparent increase in metacognitive sensitivity may trivially be explained by an increase in task performance if one of these uncorrected measures is employed.
@@
In contrast, metacognitive calibration refers to whether an observer reports a generally appropriate level of confidence given their probability of being correct. For example, if an individual—or an LLM—reports 75% confidence across multiple trials, calibration can be considered optimal when the actual proportion of correct answers in those trials is also 75% (Maniscalco et al., 2024). The Expected Calibration Error (ECE) is often used in computer science research to summarize the overall discrepancy between confidence and accuracy. ECE is typically computed by binning predictions according to confidence levels and comparing average confidence within each bin to the empirical accuracy. Calibration curves—graphs plotting model confidence against observed accuracy—are also commonly used to visualize calibration performance (Figure 1, bottom row). A perfectly calibrated system would exhibit a calibration line that falls on the diagonal (i.e., predicted confidence equals actual accuracy at all levels). Deviations from this line reflect systematic biases such as overconfidence (when predicted confidence exceeds accuracy) or underconfidence (when accuracy exceeds confidence). However, note that in the literature on human metacognition, apparent over- or under-confidence may in fact be mathematically optimal when considering reward functions or the observer’s global strategy or goals, such as whether it is more desirable to maximally avoid high-confidence errors given the consequences of such errors in the environment (Maniscalco et al., 2024).
@@
Figure 1: Demonstrations of confidence-accuracy relationships using cartoons and an empirical example based on results from GPT-3.5 (Steyvers et al., 2025b) and a confidence finetuned GPT-4.1-mini model (Steyvers et al., 2025a), focusing on metacognitive sensitivity and calibration. Top row: Confidence distributions for correct (green) and incorrect (blue) answers allow assessment of metacognitive sensitivity. Illustrative results show examples of different degrees of separations between the distributions reflecting different degrees of metacognitive sensitivity. The empirical results using GPT-3.5 and GPT-4.1-mini show modest separation, with the area under the curve (AUC = 0.778 and AUC=0.83) reflecting the probability that a randomly selected correct answer is assigned higher confidence than a randomly selected incorrect answer. Bottom row: Metacognitive calibration can be seen by plotting accuracy as a function of confidence. The illustrative results show examples of over-confidence, under-confidence, and properly calibrated confidence (points directly along the diagonal). The GPT-3.5 results, based on implicit confidence signals from token likelihoods on multiple-choice questions (MMLU), show overconfidence—predicted confidence exceeds actual accuracy. In contrast, GPT-4.1-mini was finetuned to generate explicit verbal confidence estimates on short-answer trivia questions (TRIVIAQA), yielding improved calibration.
@@ 4 Comparing Human and LLM Metacognitive Architecture and Behavior
There are several notable parallels between how humans and LLMs not only generate and calibrate confidence, but also express it (see Table 1 for an overview). These similarities may seem surprising, given the fundamental architectural and cognitive differences between humans and LLMs, yet important differences also remain; exploring these differences and their consequences on collaborative behavior may be key to effective human-LLM collaboration.
@@ 4.1 Similarities Between Humans and LLMs
One point of convergence may be with some mechanisms thought to generate confidence. In LLMs, one approach to estimating confidence leverages their probabilistic nature: the model can be prompted multiple times with the same question, and confidence can be inferred from the consistency of responses (Xiong et al., 2024; Liu et al., 2025) — similar to the other implicit measures of metacognition discussed above. Interestingly, this approach is similar to a proposed theoretical framework for human confidence, in which subjective certainty arises from the self-consistency of internally generated candidate answers (Koriat, 2012). Although developed independently in AI and cognitive psychology, both approaches suggest that consistency across internally simulated alternatives may serve as a basis for confidence.
Another similarity concerns the outwardly-visible behavioral patterns of calibration and sensitivity. Recent work has shown that when given the same task, LLMs and humans both tend to exhibit overconfidence, and both can achieve a similar degree of metacognitive sensitivity—that is, their confidence ratings are similarly diagnostic of accuracy (Cash et al., 2025). (Note, however, that this study used AUROC2—which is confounded with accuracy (Fleming & Lau, 2014)—to quantify metacognitive sensitivity, but did not control for accuracy across the LLMs and humans.) The tendency toward overconfidence has long been observed in human cognition (Kelly & Mandel, 2024), and appears to extend to large language models as well, possibly due to inductive biases or training data characteristics (Zhou et al., 2024).
@@
Finally, metacognition in humans is thought to rely on introspection-like processes, defined specifically by the privileged access we have to our own thoughts over those of others (i.e., the difference between metacognition and theory of mind). Similarly, it has been suggested that LLMs can better predict their own behavior than the behavior of another LLM, which some researchers interpret to imply the presence of such privileged access in the LLMs tested (Binder et al., 2025). Evidence of introspective-like capacities may also come from LLMs’ demonstrated ability to describe their own behaviors after training even when those behaviors are not explicitly described in their training data (such as preferring risky choices), including behaviors displayed via “backdoors” in which models show unexpected or undesirable behaviors under certain trigger conditions (such as holding a goal to elicit certain behaviors from a human user) (Betley et al., 2025). In that study, the researchers asked the models to describe their ‘tendencies’ or ‘goals’ in general, separate from a specifically prompted behavior, and found that they could describe these predilections or goals accurately—suggesting some degree of introspective access that they can explicitly report.
Table 1: Comparison of human and LLM metacognitive capabilities.
Capability Humans LLMs
Expressing confidence Flexibly and automatically report confidence across many domains; humans often appear to exhibit overconfidence (Kelly & Mandel, 2024), but this may reflect strategic tradeoffs (Maniscalco et al., 2024) Default models have limited capacity to report calibrated numeric confidence that discriminates between correct and incorrect answers; tend to be overconfident when expressing confidence verbally or numerically (Steyvers et al., 2025b; Zhou et al., 2024); finetuning can improve both sensitivity and calibration (Steyvers et al., 2025a)
Mechanisms for assessing uncertainty Confidence may reflect internal consistency or access to task-relevant information (Koriat, 2012), or formation of second-order beliefs (Peters, 2022) Token likelihoods and response consistency are used to estimate uncertainty (Kadavath et al., 2022; Liu et al., 2025)
Metacognitive training Some evidence for improvement with training, mostly in calibration; no evidence for gains in metacognitive sensitivity (Haddara & Rahnev, 2022; Kelly & Mandel, 2024; Rouy et al., 2022) Finetuning on metacognitive tasks can improve confidence calibration and sensitivity but any gains in metacognitive sensitivity show only partial generalization to other domains (Steyvers et al., 2025a; Stengel-Eskin et al., 2024)
Metacognitive control Ability to self-direct learning and offload cognition strategically (Gureckis & Markant, 2012; Gilbert, 2024) Ability to integrate external tools (e.g., search engines, calculators) enabling a form of cognitive offloading
Introspection Privileged introspective access to at least some internal processes Limited introspective-like behaviors, such as predicting their outputs better than others (Binder et al., 2025; Betley et al., 2025)